# Mitigating the Modality Gap: Few-Shot Out-of-Distribution Detection with Multi-modal Prototypes and Image Bias Estimation

### v1

@@ Abstract
Existing vision-language model (VLM)-based methods for out-of-distribution (OOD) detection typically rely on similarity scores between input images and in-distribution (ID) text prototypes. However, the modality gap between image and text often results in high false positive rates, as OOD samples can exhibit high similarity to ID text prototypes. To mitigate the impact of this modality gap, we propose incorporating ID image prototypes along with ID text prototypes. We present theoretical analysis and empirical evidence indicating that this approach enhances VLM-based OOD detection performance without any additional training. To further reduce the gap between image and text, we introduce a novel few-shot tuning framework, suPreMe, comprising biased prompts generation (BPG) and image-text consistency (ITC) modules. BPG enhances image-text fusion and improves generalization by conditioning ID text prototypes on the Gaussian-based estimated image domain bias; ITC reduces the modality gap by minimizing intra- and inter-modal distances. Moreover, inspired by our theoretical and empirical findings, we introduce a novel OOD score $S_{\textit{GMP}}$, leveraging uni- and cross-modal similarities. Finally, we present extensive experiments to demonstrate that suPreMe consistently outperforms existing VLM-based OOD detection methods.
@@ 1 Introduction
Figure 1: Standard VLM-based OOD detection methods [45, 46, 36] only utilize ID text prototypes ($\Diamond$) for identifying OOD samples. In comparison, suPreMe employs ID image prototypes ($\square$) to complement ID text prototypes, reducing the impact of the image-text modality gap [39] and sharpening the boundary between ID and OOD data. “Img.” and “Emb.” represent image and embedding.
Detecting out-of-distribution (OOD) samples [18, 38, 41, 21, 12, 60, 56] is essential for the real-world deployment of machine learning models [16, 53, 65], as novel samples may emerge and should be flagged for careful consideration. Recently, inspired by the power of vision-language foundation models (VLMs) [53, 14, 20], novel approaches to OOD detection using VLMs [12, 11, 45, 1, 40, 46, 51, 61, 2, 4, 24, 36, 69] have gained significant attention. Early VLM-based OOD detection works [45, 61, 24, 4] mainly focus on using CLIP [53] in a zero-shot setting, where only the in-distribution (ID) class names are utilized. For example, Maximum Concept Matching (MCM) [45] measures the similarity between input images and text embeddings of ID classes, also referred to as ID text prototypes, in the joint vision-language representation space of CLIP. MCM then uses this similarity score to differentiate OOD and ID samples, on the basis that ID samples should have higher similarity scores.
Following MCM, other zero-shot methods [61, 24, 4] aim at generating OOD text prototypes by querying large language models (LLMs) [26, 49] or WordNet [44]. These methods obtain OOD scores by evaluating the similarity difference between ID and OOD text prototypes. To further improve performance, various few-shot tuning OOD detection methods [40, 36, 69, 2, 48] have been proposed. These methods mainly target learning OOD (negative) text prototypes. For example, NegPrompt [36] learns OOD text prototypes by minimizing the similarity between OOD text prototypes and ID data with contrastive learning.
While these approaches have shown promise using CLIP, they use only text prototypes and thus are limited by the modality gap [39] between image and text modalities. Liang et al. [39] demonstrate that CLIP’s image and text embeddings are located in two clearly separate regions of the embedding space. As a result, OOD images may also exhibit high similarity to ID text prototypes, since high similarity can stem from either true semantic similarity or mere spatial closeness to the ID text prototypes, causing a risk of increased false positives. This observation raises a critical question about the impact of the modality gap on VLM-based OOD detection, leading to our first research question:
@@
Given these findings, i.e., that the modality gap significantly affects VLM-based OOD detection, we conclude that it is necessary to explore approaches mitigating the gap to improve reliability. This leads us to our second research question:
@@
To address RQ2, we propose a novel few-shot tuning, multi-modal prototype-based method for OOD detection with CLIP, termed suPreMe. suPreMe comprises biased prompts generation (BPG) and image-text consistency (ITC) modules. BPG introduces a Gaussian-based image domain bias to enhance image-text fusion and improve generalization. ITC reduces the modality gap by minimizing inter- and intra-modal distances. suPreMe also makes use of a new OOD score, which we call the Generalized Multi-modal Prototype OOD score, denoted $S_{\textit{GMP}}$.
BPG. To improve multi-modal fusion and generalization, in contrast to previous OOD detection methods [36, 46, 2] that only utilize learnable contexts for generating text prototypes, BPG conditions text prototypes on three components: learnable contexts, the image embedding, and the ID image domain bias. The image embedding and ID image domain bias facilitate image-text fusion, enhancing cross-modal alignment. The third component, the Gaussian-based ID image domain bias, captures the distribution of ID image embeddings to improve generalization.
ITC. To minimize the modality gap, we first map the image embedding $I$ to the text domain as $I^{{}^{\prime}}=f_{i2t}(I)$ with the image-to-text mapping $f_{i2t}(\cdot)$. To ensure it aligns with ID text prototypes, we introduce the inter-modal loss $\ell_{\textit{inter}}$. Additionally, to avoid information loss, the intra-modal loss $\ell_{\textit{intra}}$ is applied between the original embedding $I$ and the reconstructed image embedding $\hat{I}=f_{t2i}(I^{{}^{\prime}})$ using the text-to-image mapping $f_{t2i}(\cdot)$.
$\boldsymbol{S_{\textit{GMP}}}$. Building on our findings from RQ1, we introduce a new OOD score, $S_{\textit{GMP}}$, which integrates uni- and cross-modal similarities. While previous methods rely solely on the similarity between ID text prototypes and input image embedding, $S_{\textit{GMP}}$ incorporates the similarity between multi-modal input embeddings—the image embedding $I$ and the mapped image embedding $f_{i2t}(I)$—and ID multi-modal (image and text) prototypes, respectively. This allows $S_{\textit{GMP}}$ to balance inter- and intra-modal similarities, enhancing robustness and performance.
@@
  • (RQ2) To further mitigate the modality gap, we propose a novel few-shot tuning framework, suPreMe, comprising biased prompts generation (BPG) and image-text consistency (ITC) modules.
@@
  • (RQ2) Extensive experiments across multiple benchmarks, including ImageNet-1k [7], ImageNet-100 [7], ImageNet-10 [7], and ImageNet-20 [7], demonstrate that suPreMe outperforms existing OOD detection methods and achieves new state-of-the-art performance. We plan to release our code upon publication.
@@ 2 Related Work
Out-of-distribution (OOD) detection. OOD detection [18, 22, 41, 65, 68] aims to discriminate ID samples and OOD samples. Traditional methods include post-hoc [55, 30, 50, 56], confidence enhancement [9, 17, 62, 64, 47], and outlier exposure methods [19, 23, 10, 28, 21, 42, 71, 12], by estimating the OOD probability using logits [18, 22, 41, 38] or feature representations [60, 55]. Recently, the field has shifted towards exploiting large-scale pre-trained VLMs [53, 14, 20], i.e., CLIP [53], to enhance OOD detection, transitioning from uni-modal to multi-modal approaches. MCM [45] employs CLIP and explores the effects of softmax and temperature scaling in a zero-shot manner. Subsequent works can be roughly categorized into zero-shot and few-shot tuning methods. Zero-shot methods [45, 24, 4, 61] primarily focus on identifying negative classes by querying LLMs [26, 49] or WordNet [44]. In parallel, few-shot tuning methods [40, 36, 69, 2, 48] also mainly target finding negative text prompts (OOD text prototypes). For example, CATEX [40], LSN [48], and NegPrompt [36] learn negative prompts by minimizing the similarity between negative prompts and ID training data.
@@
Prompt tuning. Prompt tuning originates in Natural Language Processing [31, 37, 54] as a method for automating template/prompt creation in models such as BERT [8] and GPT [49]. For example, AutoPrompt [54] is a gradient-based approach for identifying “optimal” prompts, replacing manually designed prompts. By adapting prompts in the input embedding space based on downstream data, prompt tuning offers a parameter-efficient alternative to fine-tuning. Recently, prompt tuning has been applied in computer vision models [27, 72, 73]. Notably, CoOp [73] and CoCoOp [72], as representative methods in visual prompt tuning (VPT), employ learnable prompts optimized by minimizing classification loss to improve CLIP’s performance.
@@ 3.1 Problem Setup
Figure 2: Overview of suPreMe. The two novel modules, i.e., BPG (biased prompts generation) and ITC (image-text consistency), are designed to minimize the modality gap. 1) During the few-shot fine-tuning stage, BPG generates image domain-biased prompts, conditioned on the estimated image domain bias and the mapped image embedding $m(I)$ for better image-text fusion. ITC minimizes the modality gap directly by the intra- and inter-modal losses ($\ell_{\textit{intra}}$ and $\ell_{\textit{inter}}$) with the text-to-image mapping $f_{t2i}(\cdot)$ and the image-to-text mapping $f_{i2t}(\cdot)$. 2) During inference, image prototypes are obtained by averaging each class’s base ID image embeddings (Eq. 3). The proposed $S_{\textit{GMP}}$ (Eq. 11) is calculated based on the maximum similarity between the multi-modal embeddings (the image embedding $I_{t}$ and mapped image embedding $I_{t}^{{}^{\prime}}$) and ID multi-modal prototypes (i.e., text prototypes $\{P_{t,c}\}_{c\in[C]}$ and image prototypes $\{P_{i,c}\}_{c\in[C]}$). $S_{\textit{MCM}}$ refers to the MCM score [45].
Let $\mathcal{X}$ and $\mathcal{Y}=\{y_{1},\ldots,y_{C}\}$ represent the feature space and ID label space, respectively, where $C$ is the number of ID classes.
@@ 3.2 Revisiting CLIP, MCM, and Prompt Tuning
CLIP [53] is a foundational VLM pre-trained on a web-scale image-text dataset using self-supervised contrastive learning [5], which achieves strong zero-shot classification results on numerous benchmark datasets [29, 32, 3, 43, 52]. Specifically, for classification, CLIP first incorporates class labels $\mathcal{Y}=\{y_{c}\}_{c\in[C]}$, e.g., “cat” and “dog”, into fixed pre-designed, rather than learned, text prompts, e.g., “a photo of a [CLASS]”. These prompts combined with class names are processed by CLIP’s text encoder $f_{\textit{text}}(\cdot)$ to generate text embeddings $\{P_{t,c}\}_{c\in[C]}$, where $P_{t,c}=f_{\textit{text}}(\texttt{a photo of a }y_{c})$. Given an image $\mathbf{x}$, the image embedding $I$ is obtained by the image encoder $f_{\textit{image}}(\cdot)$ as $I=f_{\textit{image}}(\mathbf{x})$. The predicted class label is $\hat{y}=\operatorname*{arg\,max}_{c\in[C]}\operatorname{cos}(I,P_{t,c})$, where $\operatorname{cos}(\cdot,\cdot)$ denotes the cosine similarity. The zero-shot classification score is calculated as $p=\operatorname*{Softmax}([\operatorname{cos}(I,P_{t,1}),\ldots,\operatorname{cos}(I,P_{t,C}])$, where $\operatorname*{Softmax}(\cdot)$ is the softmax function.
@@
$S_{\textit{MCM}}(I,\{P_{t,c}\}_{c\in[C]})=\max_{c\in[C]}\frac{\exp(\operatorname{cos}(I,P_{t,c})/\tau)}{\sum_{j\in[C]}\exp(\operatorname{cos}(I,P_{t,j})/\tau)},$ (1)
where $\tau$ is the temperature, and $\exp(\cdot)$ is the exponential function.
Visual prompt tuning [73, 72, 15, 67]. To improve CLIP’s performance when target training data is accessible, CoOp [73] replaces manually designed prompt templates with learnable (soft) prompts as $\textit{CoOp}_{c}=[V_{1},\ldots,V_{L},y_{c}]$, where $c\in[C]$, $L$ is the length of learnable prompts, and each $V_{*}$ represents a learnable vector. The class-wise text embedding is then generated as $P^{\textit{CoOp}}_{t,c}=f_{\textit{text}}(\textit{CoOp}_{c}),\forall c\in[C]$. With CLIP’s text and image encoder parameters frozen, the learnable prompts are optimized using the cross-entropy loss,
$\ell_{\textit{ID}}=-\log\frac{\exp({\operatorname{cos}(I,P_{t,y}^{\textit{CoOp}}})/\tau)}{\sum_{c\in[C]}\exp({\operatorname{cos}(I,P_{t,c}^{\textit{CoOp}}})/\tau)}\,,$ (2)
where $y$ is the ground truth label for the input image. Following CoOp, to avoid potential overfitting on the training data, CoCoOp [72] conditions prompts on the image embedding $I$ as $\textit{CoCoOp}_{c}=[V_{1}+m(I),\ldots,V_{L}+m(I),y_{c}]$, where $m(\cdot)$ is a two-layer multi-layer perception (MLP, Linear-ReLU-Linear).
@@ 4 Methodology
In this section, we first answer RQ1 by presenting our hypothesis and findings in Sec. 4.1, where we demonstrate that the modality gap negatively impacts performance and that incorporating image prototypes alleviates this effect. Then, to mitigate the modality gap (RQ2), we introduce suPreMe in Sec. 4.2 with two modules and a novel OOD score, $S_{\textit{GMP}}$, as shown in Fig. 2. Note that in this section, we denote the input image embedding by $I_{t}$.

### v2

@@ Abstract
Existing vision-language model (VLM)-based methods for out-of-distribution (OOD) detection typically rely on similarity scores between input images and in-distribution (ID) text prototypes. However, the modality gap between image and text often results in high false positive rates, as OOD samples can exhibit high similarity to ID text prototypes. To mitigate the impact of this modality gap, we propose incorporating ID image prototypes along with ID text prototypes. We present theoretical and empirical evidence indicating that this approach enhances VLM-based OOD detection performance without any additional training. To further reduce the gap between image and text, we introduce a novel few-shot tuning framework, suPreMe, comprising biased prompt generation (BPG) and image-text consistency (ITC) modules. BPG enhances image-text fusion and improves generalization (prevents overfitting on the training data) by conditioning ID text prototypes on the Gaussian-based estimated image domain bias; ITC reduces the modality gap by minimizing intra- and inter-modal distances. Moreover, inspired by our theoretical and empirical findings, we introduce a novel OOD score $S_{\textit{GMP}}$, leveraging uni- and cross-modal similarities. Finally, extensive experiments demonstrate that suPreMe consistently outperforms existing VLM-based OOD detection methods.
@@ 1 Introduction
[…]
(a) Previous methods
(b) Our method: suPreMe
[…]
Detecting out-of-distribution (OOD) samples [18, 38, 41, 21, 11, 58, 54] is essential for the real-world deployment of machine learning models [16, 51, 63], as novel samples may emerge and should be flagged for careful consideration. Recently, inspired by the power of vision-language foundation models (VLMs) [51, 14, 20], novel approaches to OOD detection using VLMs [11, 10, 44, 1, 40, 45, 50, 59, 2, 3, 24, 36, 70, 69, 35] have gained significant attention. Early VLM-based OOD detection works [44, 59, 24, 3] mainly focus on using CLIP [51] in a zero-shot setting, where only the in-distribution (ID) class names are utilized. For example, Maximum Concept Matching (MCM) [44] measures the similarity between input images and text embeddings of ID classes, also referred to as ID text prototypes, in the joint vision-language representation space of CLIP. MCM then uses this similarity score to differentiate OOD and ID samples, on the basis that ID samples should have higher similarity scores.
Following MCM, other zero-shot methods [59, 24, 3] aim at generating OOD (negative) text prototypes by querying large language models (LLMs) [26, 48] or WordNet [43]. These methods obtain OOD scores by evaluating the similarity difference between ID and OOD text prototypes. To further improve performance, various few-shot tuning OOD detection methods [40, 36, 70, 2, 47] have been proposed. These methods mainly target learning OOD (negative) text prototypes. For example, NegPrompt [36] learns OOD text prototypes by minimizing the similarity between OOD text prototypes and ID data with contrastive learning.
Figure 2: Examples where OOD images (iNaturalist) are located closer to ID text prototypes compared to ID images (ImageNet-1k) generated by CLIP (ViT-B/16). We also visualize the embeddings generated by our suPreMe in Fig. 5, which has smaller gaps between image and text embeddings.
While these approaches have shown promise using CLIP, they use only text prototypes and thus are limited by the modality gap [39] between image and text modalities. Liang et al. [39] demonstrate that CLIP’s image and text embeddings are located in two clearly separate regions of the embedding space. As a result, OOD images may also exhibit high similarity to ID text prototypes (Fig. 2, more examples are in Fig. 6), since high similarity can stem from either true semantic similarity or mere spatial closeness to the ID text prototypes, causing a risk of increased false positives. This observation raises a critical question about the impact of the modality gap on VLM-based OOD detection, leading to our first research question:
@@
Given these findings, i.e., that the modality gap significantly affects VLM-based OOD detection, we conclude that it is necessary to explore approaches mitigating the gap to improve reliability. This leads us to our second question:
@@
To address RQ2, we propose a novel few-shot tuning, multi-modal prototype-based method for OOD detection with CLIP, termed suPreMe. suPreMe comprises biased prompt generation (BPG) and image-text consistency (ITC) modules. BPG introduces a Gaussian-based image domain bias to enhance image-text fusion and generalization on ID data unseen during fine-tuning by preventing overfitting on the limited training ID data. ITC reduces the modality gap by minimizing inter- and intra-modal distances as CLIP remains frozen. suPreMe also makes use of a new OOD score, which we call the Generalized Multi-modal Prototype OOD score, denoted $S_{\textit{GMP}}$.
BPG. In contrast to previous OOD detection methods that only utilize learnable contexts for text prototypes [36, 45, 2], BPG conditions text prototypes on three components: learnable contexts, the image embedding, and the ID image domain bias. The image embedding facilitates image-text fusion, enhancing cross-modal alignment. The third component, the Gaussian-based ID image domain bias, captures the distribution of ID image embeddings to improve generalization on unseen ID data and avoid overfitting.
ITC. As the CLIP remains frozen during training, to minimize the modality gap, we introduce intra- and inter- modal losses with two mappings. First, the image embedding $I$ is mapped to the text domain as $I^{\prime}=f_{\textit{img-txt}}(I)$ with the image-to-text mapping $f_{\textit{img-txt}}$. To ensure it aligns with ID text prototypes, we employ the inter-modal loss $\ell_{\textit{inter}}$. Additionally, to avoid information loss, the intra-modal loss $\ell_{\textit{intra}}$ is applied between the original embedding $I$ and the reconstructed image embedding $\hat{I}=f_{\textit{txt-img}}(I^{\prime})$ using the text-to-image mapping $f_{\textit{txt-img}}$.
$\boldsymbol{S_{\textit{GMP}}}$. Building on our findings, we introduce a new OOD score, $S_{\textit{GMP}}$, which integrates uni- and cross-modal similarities. While previous methods rely solely on the similarity between ID text prototypes and input image embedding, $S_{\textit{GMP}}$ incorporates the similarity between multi-modal input embeddings—the image embedding $I$ and the mapped image embedding $f_{\text{img-txt}}(I)$—and ID multi-modal (image and text) prototypes, respectively.
@@
  • (RQ2) To further mitigate the modality gap, we propose a novel few-shot tuning framework, suPreMe, comprising biased prompt generation (BPG) and image-text consistency (ITC) modules.
@@
  • (RQ2) Extensive experiments across multiple benchmarks [6, 68] demonstrate that suPreMe outperforms existing OOD detection methods.
@@ 2 Related Work
Out-of-distribution (OOD) detection. OOD detection [18, 22, 41, 63, 68] aims to discriminate ID samples and OOD samples. Traditional methods include post-hoc [53, 30, 49, 54], confidence enhancement [8, 17, 60, 62, 46], and outlier exposure methods [19, 23, 9, 28, 21, 42, 72, 11]. Recently, the field has shifted towards exploiting large-scale pre-trained VLMs [51, 14, 20], i.e., CLIP [51], to enhance OOD detection. MCM [44] employs CLIP and explores the effects of softmax and temperature scaling in a zero-shot manner. Subsequent works can be roughly categorized into zero-shot and few-shot tuning methods. Zero-shot methods [44, 24, 3, 59] primarily focus on identifying negative classes by querying LLMs [26, 48] or WordNet [43]. In parallel, few-shot tuning methods [40, 36, 70, 2, 47] also mainly target finding negative text prompts (OOD text prototypes). For example, CATEX [40], LSN [47], and NegPrompt [36] learn negative prompts by minimizing the similarity between negative prompts and ID training data. Following those works, SynOOD [35] generates OOD images using image generative models and LLMs to finetune the image and text encoders. On the other side, AdaNeg [69] introduces a memory bank during testing to selectively cache discriminative features from test images, representing the targeted OOD distribution. The memory bank further helps in distinguishing new OOD images.
@@
Prompt tuning. Prompt tuning originates in Natural Language Processing [31, 37, 52] as a method for automating template/prompt creation in models such as BERT [7] and GPT [48]. For example, AutoPrompt [52] is a gradient-based approach for identifying “optimal” prompts, replacing manually designed prompts. Recently, prompt tuning has been applied in computer vision models [27, 73, 74]. Notably, CoOp [74] and CoCoOp [73], as representative methods in visual prompt tuning (VPT), employ learnable prompts optimized by minimizing classification loss to improve CLIP’s performance.
@@ 3.1 Problem Setup
Figure 3: Overview of suPreMe. The two novel modules, i.e., BPG (biased prompt generation) and ITC (image-text consistency), are designed to minimize the modality gap. (a) During the few-shot fine-tuning stage, (b) BPG generates image domain-biased prompts, conditioned on the estimated image domain bias and the mapped image embedding $m(I)$ for better image-text fusion. (c) ITC minimizes the modality gap directly by the intra- and inter-modal losses ($\ell_{\textit{intra}}$ and $\ell_{\textit{inter}}$) with the text-to-image mapping $f_{\textit{txt-img}}$ and the image-to-text mapping $f_{\textit{img-txt}}$. (d) During inference, image prototypes are obtained by averaging each class’s base ID image embeddings (Eq. 3). The proposed $S_{\textit{GMP}}$ (Eq. 10) is calculated based on the maximum similarity between the multi-modal embeddings (the image embedding $I$ and mapped image embedding $I^{\prime}$) and ID multi-modal prototypes (i.e., text prototypes $\{P_{\textit{txt},c}\}_{c=1}^{C}$ and image prototypes $\{P_{\textit{img},c}\}_{c=1}^{C}$). $S_{\textit{MCM}}$ refers to the MCM score [44].
Let and $\mathcal{Y}=\{y_{1},\ldots,y_{C}\}$ represent the set of ID class labels, where $C$ is the number of ID classes.
@@ 3.2 Revisiting CLIP, MCM, and Prompt Tuning
CLIP [51] is a foundational VLM pre-trained on a web-scale image-text dataset using self-supervised contrastive learning [4]. Specifically, for classification, CLIP first incorporates class labels $\mathcal{Y}=\{y_{c}\}_{c=1}^{C}$, e.g., “cat”, into fixed pre-designed, rather than learned, text prompts, e.g., “a photo of a label ”. These prompts combined with class labels are processed by CLIP’s text encoder $f_{\textit{text}}(\cdot)$ to generate text embeddings $\{P_{\textit{txt},c}\}_{c=1}^{C}$, where $P_{\textit{txt},c}=f_{\textit{text}}(\text{``}\texttt{a photo of a }\{y_{c}\}\text{''})$. Given an image $\mathbf{x}$, the image embedding $I$ is obtained by the image encoder $f_{\textit{image}}(\cdot)$ as $I=f_{\textit{image}}(\mathbf{x})$. The index of the predicted class label is given by $\operatorname*{arg\,max}_{c\in\{1..C\}}\cos(I,P_{\textit{txt},c})$, where $\cos(\cdot,\cdot)$ denotes the cosine similarity.
@@
$S_{\textit{MCM}}(I,\{P_{\textit{txt},c}\}_{c=1}^{C})=\\ \max_{c\in\{1..C\}}\frac{\exp(\cos(I,P_{\textit{txt},c})/\tau)}{\sum_{j=1}^{C}\exp(\cos(I,P_{\textit{txt},j})/\tau)},$ (1)
where $\tau$ is the temperature and $\exp(\cdot)$ is the exponential function.
Visual prompt tuning [74, 73, 15, 65]. To improve CLIP’s performance when target training data is accessible, CoOp [74] replaces manually designed prompt templates with learnable soft prompts (context tokens) $\textit{CoOp}_{c}=[V_{1},\ldots,V_{L},y_{c}]$ for class index $c$. $L$ is the length of learnable prompts and each $V_{i}$ represents a learnable vector. The class-wise text embedding is then generated as $P^{\textit{CoOp}}_{\textit{txt},c}=f_{\textit{text}}(\textit{CoOp}_{c})$ for class index $c$. With CLIP’s text and image encoder parameters frozen, the learnable prompts are optimized using the cross-entropy loss:
$\ell_{\textit{ID}}(I)=-\log\frac{\exp({\cos(I,P_{\textit{txt},c}^{\textit{CoOp}}})/\tau)}{\sum_{j=1}^{C}\exp({\cos(I,P_{\textit{txt},j}^{\textit{CoOp}}})/\tau)}\,$, (2)
where $c$ is the index of the ground truth label for the input image. Following CoOp, to avoid potential overfitting on the training data, CoCoOp [73] conditions prompts on the image embedding $I$ as $\textit{CoCoOp}_{c}=[V_{1}+m(I),\ldots,V_{L}+m(I),y_{c}]$, where $m(\cdot)$ is a two-layer multi-layer perception (MLP, Linear-ReLU-Linear).
@@ 4 Methodology
In this section, we first answer RQ1 by presenting our hypothesis and findings in Sec. 4.1, where we demonstrate that the modality gap negatively impacts performance and that incorporating image prototypes alleviates this effect. Then, to mitigate the modality gap (RQ2), we introduce suPreMe in Sec. 4.2 with two modules and a novel OOD score, $S_{\textit{GMP}}$. We provide an overview of suPreMe in Fig. 3.